Two electrons have been removed from each atom.

Find the distance between two such atoms, if they

repel each other with a force of 8.8´10-9N, when placed

in free space.


1
Expert's answer
2021-04-15T20:40:49-0400

Answer

Distance us given by

"R=\sqrt{\frac {ke^2}{F}}\\=\sqrt{\frac {9\times10^9\times1.6\times1.6\times10^{-38}}{8.8\times10^{-9}}}\\=1.62\times10^{-10}m"